OpenStudy (anonymous):

Suppose that \( \large p \) is a prime, then \[ \large 2^{p-1}\left(2^p -1 \right) \] ends with a 6 or 8.

OpenStudy (anonymous):

let's consider \(2^{p-1}(2^p-1)\mod10\)... there is a clear pattern \(2^n\mod 10\):$$2^0\equiv1\mod10\\2^1\equiv2\mod10\\2^2\equiv4\mod10\\2^3\equiv8\mod10\\2^4\equiv6\mod10\\2^5\equiv2\mod10\\\dots$$so clearly for odd \(n\) we see that \(2^n\) cycles \(2,8,2,8,2\dots\). either \(2^p\equiv2\) or \(2^p\equiv8\)

OpenStudy (anonymous):

when \(p = 2\), its clearly true. we only have to prove it for odd primes. and we can use fermat theorem i guess ?

OpenStudy (anonymous):

if \(2^p\equiv2\) then either \(2^{p-1}\equiv1\) (for \(p=1\)) or \(2^{p-1}\equiv6\). since \(p\) is prime we know then that \(2^{p-1}\equiv6\) ergo \(2^p-1\equiv1\) and \(2^{p-1}(2^p-1)\equiv6\cdot1\equiv6\)

OpenStudy (anonymous):

if \(2^p\equiv8\) then we know \(2^{p-1}\equiv4\) ergo \(2^p-1\equiv7\) so \(2^{p-1}(2^p-1)\equiv4\cdot7\equiv8\) Q.E.D.

OpenStudy (anonymous):

yep, that'll do it :-)

OpenStudy (anonymous):

i understood \(2^p \equiv 8\) case comletely however i dont get the \(2^p\equiv 2 \) case yet :|

OpenStudy (anonymous):

when \(2^p\equiv2 \mod 10\), \(2^{p-1} \equiv 1 \mod 10\) ... how can we divide like this ? 10 is not a prime right

OpenStudy (anonymous):

I don't divide (clearly \(2\) is non-invertible) but I just look at the pattern

OpenStudy (anonymous):

oh your first replyy...

OpenStudy (anonymous):

you can show very easily that \(2^n\) is periodic for \(n>0\), the only "odd" term is \(n=0\

OpenStudy (anonymous):

but yes

OpenStudy (anonymous):

so the whole proof is based on cyclicity... we look at what comes before the powers for remainders 2 and 8. got it :) and it looks great, thank you :)

OpenStudy (anonymous):

NP and yeah it's all about how the powers are periodic

OpenStudy (anonymous):

Here is a prrof that I know: If p=2, then \( 2^{p-1} \left( 2^p-1 \right)=6\) if p>2, then p=4m +1 or p= 4m +3 othereise, p is divisible by 2 Let us consider the first case p=4m +1 \[ 2^{p-1} \left( 2^p-1 \right)= 2^{4m} \left( 2^{4m+1}-1 \right)=\\ = 16^m \left( 2\, 16^m-1 \right) \] You can show by induction that \( 16^m \) always ends up in 6. so \[ \left( 2\, 16^m-1 \right) \] ends up 1, So we are done in this case. If p=4m +3 \[ 2^{p-1} \left( 2^p-1 \right)= 2^{4m+2} \left( 2^{4m+3}-1 \right)=\\ = 4\,16^m \left( 8 \, 16^m-1 \right) \] Now \( 4 \, 16^m \) ends in 4. \( 8 \, 16^m -1 \) ends up in 7 So our number ends up in 8

OpenStudy (anonymous):

yuck! that looks a lot harder. you can make my part about periodicity rigorous very easily... and it's far less a hassle... :-p